OpenStudy (anonymous):

how to solve this word problem: 5200 tickets were sold and the receipts totaled $29,961. the adult admission is $7.50 and the child admission is $4.50. How many adult tickets were sold?

OpenStudy (nikato):

let a+ # of adult tickets sold and c+# of child tickets sold

OpenStudy (nikato):

so we have two equations tickets: a+c=5200 and the money: 7.5a +4.5c = 29961

OpenStudy (nikato):

and you can subtract a on boths dies for the first equation and get c=5200-a substitute thsi into the second equation 7.5a + 4.5(5200-a) = 29961 and now solve for a

OpenStudy (anonymous):

would the answer be 547 adult tickets sold?

OpenStudy (nikato):

no.

OpenStudy (nikato):

7.5a + 4.5(5200-a) = 29961 7.5a + 23400 -4.5a=29961 -23400 -23400 -------------------------------- 3a= 6561 ---- ----- 3 3 a=?

OpenStudy (anonymous):

a=2187?

OpenStudy (nikato):

correct!

OpenStudy (anonymous):

thank you!

OpenStudy (nikato):

no problem